OpenStudy (anonymous):

Yesterday, Asher ran 2 miles in nineteen minutes. Today, he ran 3 miles in thirty minutes. The function that describes the relationship between the number of miles he ran and the time he ran is a linear function. How long will he run tomorrow if he runs 5 miles? Explain.

OpenStudy (anonymous):



OpenStudy (anonymous):

Now that you have slope, use the slope equation: y = mx + b. y = mx + b y = 11x + b Now plug in x and y . I'm choosing (3, 30) 30 = 11(3) + b 30 = 33 + b b = -3 Now that you have b , plug back to original equation: y = 11x - 3 The problem is asking for 5 miles, so x = 5 miles y = 11(5) - 3 y = 55 - 3 y = 52 minutes. She will run 5 miles in 52 minutes the next day.

OpenStudy (anonymous):

thx